OpenStudy (anonymous):

HELPPP What is the vertex in the quadratic function y = - 4x^2 + 8x - 3 ? Write your answer as an ordered pair.

OpenStudy (johnweldon1993):

You have a quadratic in the form y = ax² + bx + c so your a = -4 b = 8 and c = -3 To find the vertex you use this formula \[\large Vertex = \frac{ -b }{ 2a }\] This will give you your x-coordinate..to find you y-coordinate....plug what you get for 'x' in the equation and solve...

OpenStudy (anonymous):

-1 ?

OpenStudy (anonymous):

is x

OpenStudy (johnweldon1993):

Close.. -b/2a so -8/2(-4) = -8/-8 =1 right?

OpenStudy (anonymous):

oh yes right.

OpenStudy (anonymous):

how would you write the equation to get y

OpenStudy (johnweldon1993):

Well now you know that at the vertex...x = 1....so to solve for 'y'...you simply plug in "1" for each 'x' in your equation. y = -4x² + 8x - 3 after replacing the 'x' with 1...we have y = -4(1)² + 8(1) - 3 y = ???

OpenStudy (anonymous):

21 ?

OpenStudy (johnweldon1993):

Not quite...remember you do exponents BEFORE you multiply... -4(1)² does not mean -4² it means -4 times 1² ....so this simply comes out to -4 so altogther we have y = -4(1)² + 8(1) - 3 y = -4 + 8 - 3 y = ?

OpenStudy (anonymous):

is this right ????

OpenStudy (anonymous):

9

OpenStudy (anonymous):

I mean 1

OpenStudy (johnweldon1993):

Right 1 would be correct...so x = 1 and y = 1...so your ordered pair is (1 , 1)

OpenStudy (johnweldon1993):

Make sense? :)

OpenStudy (anonymous):

yessssssss thanksss (:

OpenStudy (johnweldon1993):

anytime! :)

OpenStudy (anonymous):

What is the vertex in the quadratic function y = x^2 + 8x + 3? Write your answer as an ordered pair. I got this one wrong and don't know what i did wrong

OpenStudy (johnweldon1993):

Show me exactly what you did... remember the equation -b/2a...what do you get for you 'x' coordinate from this?

OpenStudy (anonymous):

-4

OpenStudy (johnweldon1993):

right....and then you plug that in for the 'x' in you equation so y = x² + 8x + 3 after plugging it in you get y = (-4)² + 8(-4) + 3 y = ???

OpenStudy (anonymous):

-13 ??

OpenStudy (johnweldon1993):

That is correct...you should not have gotten this wrong if what you got was (-4 , -13)

OpenStudy (anonymous):

first time i did it i got -45, and i was like that can't be right, its cause I didn't put in the parantheses

OpenStudy (johnweldon1993):

Oh okay....yeah just use parenthesis..it makes it easier to follow and you will be less prone to mistakes that way :)

OpenStudy (anonymous):

yeah thanks :) so muchh!

OpenStudy (johnweldon1993):

like I said...anytime lol :)